OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of one half. y - 4 = one half(x - 1) y + 4 = 2(x - 1) y + 4 = one half(x - 1) y - 4 = 2(x + 1)

OpenStudy (johnweldon1993):

remember that point slope formula i showed you? y - y1 = m(x - x1) you know your slope is m and your point is 1, -4 so y + 4 = 1/2(x-1)